OpenStudy (anonymous):

Victor calculates the percentage interest rate he needs to make five hundred dollars in simple annual interest over 5 years on a deposit of one thousand dollars. Where did he go wrong? I equals p times r times t; one thousand equals five hundred times r times five; r equals zero point four; r equals forty percent

OpenStudy (anonymous):

please help

OpenStudy (texaschic101):

I = PRT I = 500 P = 1000 T = 5 R = x now we sub 500 = (1000)(5)(t) 500 = 5000t 500/5000 = t 0.1 = t (turn to percent by multiplying by 100) 10% = t check.. I = PRT 500 = (1000)(5)(0.1) 500 = 5000(0.1) 500 = 500 (correct) ============= you just got your interest and your principal mixed up.

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)